OpenStudy (anonymous):

The point (5, -1) is a solution of the equation: y = 2x − 11 True or False

OpenStudy (anonymous):

Plug -1 into y and 5 into x in the equation, what do you get?

OpenStudy (anonymous):

just substitute tht points in equation if left hand side equal to right hand side then tht will be ur solution ..

OpenStudy (anonymous):

Just put (x,y)=(5,-1)into the given question equation I. The place of x and y

OpenStudy (anonymous):

So the equation becomes y=-1 5 So when u will get the equal terms on both sides then it will call true but if not then it will be false.

OpenStudy (anonymous):

*x=5

OpenStudy (anonymous):

its true

OpenStudy (anonymous):

Yes u r right.

OpenStudy (anonymous):

can u answer one last one?

OpenStudy (anonymous):

Yes would any body give me medal

OpenStudy (anonymous):

Calculate the slope of the points (-2, -1) and (3, 4). Answer a) 1 b) 2 c) 3 d)4

OpenStudy (anonymous):

y2-y1/x2-x1 = slope, or tangent line formula.

OpenStudy (anonymous):

M=y2-y1/x2-x1 (x1,y1)=(-2,-1) (x2,y2)=(3,4) Now just put these value in the above formula and vet the slope

OpenStudy (anonymous):

would it be 4?

OpenStudy (anonymous):

No it would be 1 Because 4-(-1)/3(-2) M=4+1/3+2 M=5/5 M=1 okay?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

U r looking cute.

OpenStudy (anonymous):

lol Ty

OpenStudy (anonymous):

Lol